OpenStudy (anonymous):

A survey of 1,000 men and women asked, "Do you earn over $50,000 per year?" The table below shows the responses for males and females: Male Female Total Income over $50,000 475 375 850 Income below $50,000 75 75 150 Total 550 450 1,000 Based on these data, are "being female" and "earning over $50,000" independent events?

OpenStudy (anonymous):

@IMStuck please help

OpenStudy (kropot72):

Let the event 'being female' be A, and let the event 'earning over $50,000' be B. Then events A and B are independent if, and only if: \[P(A \cap B)=P(A) \times P(B)\] P(A) = 450/1000 = 0.45 P(B) = 850/1000 = 0.85 \[P(A) \times P(B)=0.45\times0.85=0.3825\]

OpenStudy (anonymous):

ok im not good with this but so far i understand

OpenStudy (kropot72):

But from the table the probability of being female and earning over $50,000 is 375/1000 = 0.375 So the criterion for independence of A and B is not met, the reason being that: P(A intersection B) does not equal P(A) * P(B).

OpenStudy (anonymous):

THANK YOU!!!!!!!!!!!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

so it is No, P(being female | the person earns over $50,000) ≠ P(being female)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

or is it Yes, P(being female | the person earns over $50,000) ≠ P(being female)

OpenStudy (kropot72):

The answer is No, "being female" and "earning over $50,000" are not independent events.

OpenStudy (anonymous):

thanks!!!

OpenStudy (kropot72):

You're welcome :) Please note that there is a different test for independence to the test that I used. Only one test is necessary.